OpenStudy (anonymous):

Change 4x - 3y = -15 to slope-intercept form and state the slope and y-intercept.

OpenStudy (baldymcgee6):

Solve for y

OpenStudy (anonymous):

Ok you have the line in the form of Ax+By=C and they want you to convert it into the slope intercept form y=mx+b because then m is the slope and b is the y intercept. Ax+By=C By=C-Ax y = C/B -(A/B)x or an example with numbers. if 10x +20y = 40 20y = 40-10x y = 40/20 -(10/20)x y= -(1/2)x +2 ====== slope intercept form Now repeat with your equation